Given the financial statements for Jones Corporation and Smith Corporation:  

JONES CORPORATION
Current Assets Liabilities
Cash $ 123,300 Accounts payable $ 176,000
Accounts receivable 183,600 Bonds payable (long term) 89,600
Inventory 56,800
Long-Term Assets Stockholders' Equity
Gross fixed assets $ 554,000 Common stock $ 150,000
Less: Accumulated depreciation 152,300 Paid-in capital 70,000
Net fixed assets* 401,700 Retained earnings 279,800
Total assets $ 765,400 Total liabilities and equity $ 765,400

    

Sales (on credit) $ 1,255,000
Cost of goods sold 678,000
Gross profit $ 577,000
Selling and administrative expense 355,000
Depreciation expense 59,400
Operating profit $ 162,600
Interest expense 11,300
Earnings before taxes $ 151,300
Tax expense 82,500
Net income $ 68,800

*Use net fixed assets in computing fixed asset turnover.
†Includes $12,700 in lease payments.
  

SMITH CORPORATION
Current Assets Liabilities
Cash $ 44,000 Accounts payable $ 83,700
Marketable securities 7,900 Bonds payable (long term) 290,000
Accounts receivable 73,700
Inventory 77,500
Long-Term Assets Stockholders' Equity
Gross fixed assets $ 557,000 Common stock $ 75,000
Less: Accumulated depreciation 251,800 Paid-in capital 30,000
Net fixed assets* 305,200 Retained earnings 29,600
Total assets $ 508,300 Total liabilities and equity $ 508,300

*Use net fixed assets in computing fixed asset turnover.

   

SMITH CORPORATION
Sales (on credit) $ 1,430,000
Cost of goods sold 919,000
Gross profit $ 511,000
Selling and administrative expense 299,000
Depreciation expense 57,500
Operating profit $ 154,500
Interest expense 30,600
Earnings before taxes $ 123,900
Tax expense 59,900
Net income $ 64,000

Includes $12,700 in lease payments.
  
a. Compute the following ratios. (Use a 360-day year. Do not round intermediate calculations. Input your profit margin, return on assets, return on equity, and debt to total assets answers as a percent rounded to 2 decimal places. Round all other answers to 2 decimal places.)


Jones Corp. Smith Corp.
Profit margin 5.48 % 4.48 %
Return on assets (investments) 8.99 % 12.60 %
Return on equity 13.77 % 47.55 %
Receivable turnover 6.84 times 19.40 times
Average collection period 52.63 days 18.56 days
Inventory turnover times times
Fixed asset turnover times times
Total asset turnover times times
Current ratio times times
Quick ratio times times
Debt to total assets % %
Times interest earned times times
Fixed charge coverage times times

I answered some of them but I'm not sure if they are correct.
Answer #1
a. Compute the following ratios. (Use a 360-day year. Do not round intermediate calculations. Input your profit margin, return on assets, return on equity, and debt to total assets answers as a percent rounded to 2 decimal places. Round all other answers to 2 decimal places.)
Jones Corp. Smith Corp.
Profit margin = Net Income/Sales 5.48% 4.48%
Return on assets (investments) = Net Income/ Total Assets 8.99% 12.59%
Return on equity = Net Income/ Total shareholder Equity 13.77% 47.55%
Receivable turnover = Sales/ Account Receivables 6.836 times 19.40 times
Average collection period = 360 days/ Receivable turnover 52.67 days 18.55 days
Inventory turnover = Sales/ Inventory 22.10 times 18.45 times
Fixed asset turnover = Sales/ Net Fixed Assets 3.12 times 4.69 times
Total asset turnover = Sales/ Total assets 1.64 times 2.81 times
Current ratio = Current Assets/ Current Liabilities 2.07 times 2.43 times
Quick ratio = (Current Assets - Inventory)/Current Liabilities 1.74 times 1.50 times
Debt to total assets = Total Liabilities/ Total Assets 34.70% 73.52%
Times interest earned = Operating Profit/ Interest expense 14.39 times 5.049 times
Fixed charge coverage =( Operating Profit + Lease)/(Lease + Interest Expenses) 7.30 times 3.86 times
a. Compute the following ratios. (Use a 360-day year. Do not round intermediate calculations. Input your profit margin, return on assets, return on equity, and debt to total assets answers as a percent rounded to 2 decimal places. Round all other answers to 2 decimal places.)
Jones Corp. Smith Corp.
Profit margin = Net Income/Sales =B22/B13 =B50/B41
Return on assets (investments) = Net Income/ Total Assets =B22/C11 =B50/C37
Return on equity = Net Income/ Total shareholder Equity =B22/SUM(F8:F10) =B50/SUM(E34:E36)
Receivable turnover = Sales/ Account Receivables =B13/C5 times =B41/C31 times
Average collection period = 360 days/ Receivable turnover =360/B60 days =360/D60 days
Inventory turnover = Sales/ Inventory =B13/C6 times =B41/C32 times
Fixed asset turnover = Sales/ Net Fixed Assets =B13/C10 times =B41/C36 times
Total asset turnover = Sales/ Total assets =B13/C11 times =B41/C37 times
Current ratio = Current Assets/ Current Liabilities =SUM(C4:C6)/F4 times =SUM(C29:C32)/E29 times
Quick ratio = (Current Assets - Inventory)/Current Liabilities =(SUM(C4:C6)-C6)/F4 times =SUM(C29:C31)/E29 times
Debt to total assets = Total Liabilities/ Total Assets =SUM(F4+F5)/C11 =SUM(E29:E30)/C37
Times interest earned = Operating Profit/ Interest expense =B18/B19 times =B46/B47 times
Fixed charge coverage =( Operating Profit + Lease)/(Lease + Interest Expenses) =(B18+12700)/(12700+B19) times =(B46+12700)/(12700+B47) times
1 Glven the financial statements for Jones Corporation and Smith Corporation: 2 JONES CORPORATION 3 Current Assets 4 Cash 5 Acoounts receivable 6 Inventory 7 Long-Term Assets 8 Gross fixed assets 9 Less: Accumulated depreciation 10 Net fixod assets 11 Total assets 12 Inoome Statement 13 Sales (on crodit) 14 Cost of goods sold 15 Gross profit 16 Selling and administrative expenset 17 Deprociation expense 18 Operating profit 19 Interest expense 20 Eamings before taxes 21 Tax expense 22 Net income 23 Use net fixed assets in computing fixed asset turnover 24 tIncludes $12,700 in lease payments. Liabilities 1,23,300 Acoounts payable 1,83,600 Bonds payable (long term) 1,76,000 89,600 56,800 Stockholders Equity Common stock Paid-in capital 1,50,000 70,000 2,79,800 7,65,400 5,54,000 1,52,300 4,01,700 Rctained camings 7,65,400 Total liabilities and 12,55,000 6,78,000 5,77,000 3,55,000 1,62,600 11,300 1,51,300 68,800

25 SMITH CORPORATION 26 Current Assets 27 Cash 28 Marketable sccurities 29 Acoounts reccivable 30 Inventory 31 Long-Term Assets 32 Gross fixed assets 33 Less: Accumulated depreciation 34 Net fixod assets 35 Total asscts 36 Use net fxed assets in computing fxed asset turnover. 37 SMITH CORPORATION 38 Sales (on crodit) 39 Cost of goods sold 40 Gross profit 41 Selling and administrative expenset 42 Depreciation expense 43 Operating profit 44 Interest expense 45 Eamings before taxes 46 Tax expense 47 Net income 48 49 tIncludes $12,700 in lease payments. Liabilities 44,000 Accounts payable 83,700 7,900 Bonds payable (long term) 73,700 77,500 2,90,000 5,57,000 2,51,800 Stockholders Equity Common stock Paid-in capital 75,000 30,000 29,600 5,08,300 ,05,200 Retained eamings 5,08,300 Total liabilities and 14,30,000 9,19,000 5,11,000 2,99,000 57,500 1,54,500 30,600 1,23,900 59,900 64,000